please help i'm stuck
the length of a rectangle is 9cm greater than it's width. write an expression to represent the perimeter. simplify. also find the area in the given problem.

Answer by stanbon(75887)

the length of a rectangle is 9cm greater than it's width. write an expression to represent the perimeter. simplify. also find the area in the given problem.
------------------------
Let width be w:
Then length = w+9
----
Perimeter = 2(length + width)
---
P = 2(w+9 + w)
P = 2(2w+9)
P = 4w+18
=============
Cheers,
Stan H.